5.7.2    A hospital administrator wishes to estimate the mean weight of babies born in her

hospital. How large a sample of birth records should be taken if she wants a 95% confidence interval that is 1 pound wide? Assume that a reasonable estimate of s is 1 pound.

What if she wanted to be 99% certain?    Compute the sample size for that as well.
Answer #1

(a)

Sample Size (n) is given by:

Given:

= 0.05

From Table, critical values of Z = 1.96

s = 1

e = 1/2 = 0.5

Substituting, we get:

So,

Answer is:

16

(b)

Sample Size (n) is given by:

Given:

= 0.01

From Table, critical values of Z = 2.576

s = 1

e = 1/2 = 0.5

Substituting, we get:

So,

Answer is:

27